Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 544
Entry time: 02 Sep 2024, 09:53
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 02.09.2024 |
|
|
General Announcement:
-
CTRL:
- Today an IT update on project database at 13:00: 15min service is not available (not able to open them). Reminder will be sent.
DA:
- Note after meeting: Nothing to report.
EEE:
- Nothing to report.
VAC:
- Pressure spike in SA1 vacuum system that tripped interlock but it was quickly recovery on Thursday.
XPD:
- Nothing to report.
DET:
- Issue with Julabo chiller at AGIPD still persists as test last week did not succeed.
ITDM:
- Nothing apart of the mentioned CTRL report.
XRO:
- Vacuum spike was caused by the one motor at M1/SA1 being loose; Incident with InGa spillage in mirror chamber as a result;
- M1 Rz motor was reconnected and mirror was put in place again but now should not be moved anymore and it is interlocked.
- On Saturday beam was found back at SPB/SFX (saved trajectory at one photon energy; for others have to be checked - please do not rely on previously saved positions).
- Beam was not yet found at FXE – M3 reflection at the M3 Pop-in still seen but not further downstream.
- In case of need please do not hesitate in contacting XRO and Antje Trapp for help.
LAS:
- SA1 PP-laser NOPA alignment starting today for delivery tomorrow as AMPHOS service is done.
UNSYS:
- Quad mover stuck in cell7/SA3; needs ZZ.
- Apple-X activities during FD week had re-alignment with beam; found issue with one CAM mover in cell#25.
SPB:
- Chiller for AGIPD: tests did not get enough temperature.
FXE:
- Nothing to report.
MID:
- All fine at SA2.
HED:
- Getting ready to take beam.
SCS:
- Nothing to report.
SQS:
- Not present.
SXP:
- Not present. |